4004 Monument
- Unit Log -

2000 ~: Module retrieved from trash bin. Remainder of unit was not found, only the digital portion which has since been incorporated into the monument. The front panel matches that of an ARC RT-485 NAV/COM avionics radio.
2003 Jun: Numerous repairs:
  • IC A1 (4201) bad: replaced with reconstruction.
  • IC C2 was absent: replaced with 74LS47.
  • One clamp diode around IC A8 (4562) replaced.
  • Load resistor at IC B4-2 added, was probably on external equipment originally.
  • LEFT-INTEGER-RIGHT opto-sensor bad: replaced with alternative type.
  • RIGHT-FRACTION-LEFT and -RIGHT opto-sensors bad: removed.
  • Opto-sensor lamps replaced with 12V@35mA, all new except right-right.
  • Lamp on LEFT-1 switch burned out: replaced with that from SCAN switch.
  • 3 displays very dim: replaced.
  • Several circuit patches on board C constructed.
2004 Mar 29: Monument / power supply unit assembled, functioning.
2006 Nov 12: IC A1 replaced again with a 4201 from M900 PROM programmer.
2006 Nov 12: RIGHT-FRACTION-RIGHT and -LEFT opto-sensors filled with 2N5780.
2006 Nov 14: RIGHT-FRACTION increments for both directions of turning of the knob. RIGHT-FRACTION-RIGHT opto-sensor always closed to GND (semi-con junction drop). Resoldering / heating fixes it temporarily. After a day it reverts to improper functioning. More heating does not fix it. Removed.
2006 Nov 14: LEFT-INTEGER-RIGHT opto-sensor previously replaced with a light-sensitive diode, replaced with 2N5777. Does not work: 2N5777 does not turn on fully, collector stays around +13V with only minor drop as knob is turned. Could be an optical alignment problem. Diode re-installed.
2006 Nov 14: Spring in SCAN switch replaced with lighter spring.

4004 Monument
2009 Mar